The sum of five consecutive even integers is −20. Find the integers. List them in order
from smallest to largest.

Answer by ikleyn(52788)

.
The integers are  x,  x+2,  x+4,  x+6  and  x+8,  where  x  is the smallest of them.

The equation to find  x  is

x + (x+2) + (x+4) + (x+6) + (x+8) = -20.

Simplify and solve:

5x + 20 = -20,

5x = -40,

x = -8.

Check.   -8 + (-6) + (-4) + (-2) + 0 = -20.

Answer.   The integers are  -8,  -6,  -4,  -2  and  0.